Firm has 2400 quantity and v.c is 2.70 new supplier reduced the new v.c is 2.30 and also added advertisement cost 3000 in cost profit 10000 profit contribution 25%of the revenue? ??
T.c
C
Tr
B.e.p
New t.c
1
Expert's answer
2018-05-15T09:28:08-0400
Q = 2400, AVC1 = 2.70, AVC2 = 2.30, advertisement cost FC = 3000, profit TP = 10000, profit contribution is 25% of the revenue.
TC = FC + AVC1*Q = 3000 + 2.7*2400 = 9480.
TR = 10000/0.25 = 40000.
P = TR/Q = 40000/2400 = 16.67.
BEP1 = FC/(P - AVC1) = 3000/(16.67 - 2.7) = 214.8 units.
New TC = FC + AVC2*Q = 3000 + 2.3*2400 = 8520.